Systems Staff Meeting Notes 4/10/00

Systems Staff Meeting Notes
April 10, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Hilton Soto
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Another order of Dells is imminent.
  2. Dell report: some problems with the mouse drivers for trackballs. Also, Greg Allen has multiple problems with his machine - however, he needs to remove his extraneous programs before it can be properly diagnosed. On another machine, McAfee 4.05 (loading from Circa) killed it - changed it to 4.03 and it worked. CCC List had a note about Norton conflicting with McAfee on Dell machines.
Networking
  1. Unsecured folder on NT2 is now offline.
  2. Internet 2 had problems on Friday. The I2 link shouldn't knock anyone off, only make it slower. Should not be a problem for anyone on campus.
  3. Tech side outage on NOTIS - 4:45 to 5:15 PM last Thursday - only able to access PUBLuis & Catalog.
  4. West wiring - the cable tray installers did not get their permits until April 3rd. Judy went with them to ensure they obtained it. The permit hold-up has put Digital about two weeks behind schedule - they can't pull the wires until the cable trays are installed.
  5. NT10 - it's ready to swap over in place of the Novell (Smathers). Will wait for a dead time (like a weekend) to try it out. Hopefully it will work and Smathers will come back up after the initial testing. Rich will notify staff at the PSIT meeting.
  6. IPX question - not figured out yet, but we'd really like to kill it off.
  7. Wireless - Resource Services requested a demo for possible use in the stacks. Will is waiting to hear back from John Humphrey regarding the conduit diagrams.
  8. Hub replacement - Bobbie has a list - after reviewing it, Bill will be sent recommendations. West is not included. Will try to finalize putting Facilities in charge of wiring by next week. Hopefully they will keep us apprised of changes.
  9. HSS - multi-user project online has been proposed. Would like to use Access. It would be limited to the Autocad program. Bill will confer with them on this matter.
  10. Journalism expansion - Bobbie went over to Journalism with Gary Cornwell. He thought there wasn't any wiring in place. There are 28 ports around the Circulation desk and 3 in Priscilla's office. May have to reconsider the proposed set up. Journalism is going to pay for a live AP news feed. If that is the case, they will need machines better suited for that purpose. Will also need public printing access. There will be an expansion of the staff there - two more FTE's may be hired. They currently have a 10M feed, but 100M is available.
  11. There's a problem with the desk in Documents - the cords on the monitors are not long enough to reach the ports without causing stress. They do not want to cut a hole in the desk & it was ordered without cable routing. Bill will check on the situation.
Software Projects Status
  1. Film log demo this week.
  2. Winston made some minor modifications to the Resource Services Marcon tool. He's trying to keep the application simple. Do not want to get into making it work elsewhere.
  3. Table modifications, interface changes were made to the Personnel database.
DP Coordinator
  1. Two students are out today. This should help with the budget.
  2. Had two tape requests last week! One was for a patron, the other required making a tape copy for ICPSR - apparently they were missing some 1960 Census data.
Calendar
  1. Debra out Wednesday afternoon.
  2. Bill out late Thursday afternoon.
  3. Bill & Suzy out Friday, April 14.
  4. Winston out April 17-19.
  5. Debra out April 18-21.
  6. Rich out April 24-28.
  7. Winston out June 19-23.
  8. Bobbie out July 10-14.
  9. Will out July 24-28.


Back to the Systems Homepage

Last updated April 17, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu